Write an equation of the line that passes through the point (-1,-4) and has slope 3. Then write the equation using function notation.
Answer by rothauserc(4718)

Use the slope - intercept form
:
y = mx + b where m is the slope and b is the y intercept
:
y = 3x + b
:
now use the point coordinates (-1, -4)
:
-4 = 3(-1) + b
:
b = -4 + 3 = -1
:
*************
f(x) = 3x - 1
*************
: